OpenStudy (anonymous):

a manager at a theater knows that 900 tickets were sold for $2300.00. There are 2 ticket prices one for the balcony and one for the first floor. The first floor tickets sold for $3.00 each and the balcony sold for $2.00 each. How many of each ticket were sold? Pleaseeee help

OpenStudy (anonymous):

solve by system of two equations let x= first floor tickets and y= balcony tickets x+y=900 and 3x+2y=2300 start solve for y in x+y=900 so substract x from both sides and you get y=900-x now sub for y in other equation 3x+2(900-x)= 2300 so then you get 3x+1800-2x=2300 then add like terms and substract 1800 both sides x+1800=2300 then x=2300-1800 x= 500 first floor tickets but x value is x+y=900 solve for y 500+y=900 y=900-500 y=400 balcony tickets

OpenStudy (anonymous):

ooohhh yeah thanks im sorry i was on Instagram lol my fault